Part A) A certain first-order reaction has a rate constant of 2.40×10−2 s−1 at 25 ∘C....

Part A)

A certain first-order reaction has a rate constant of 2.40×10−2 s−1 at 25 ∘C. What is the value of k at 60 ∘C if Ea = 74.0 kJ/mol ?

Express your answer using two significant figures.

Part B)

Another first-order reaction also has a rate constant of 2.40×10−2 s−1 at 25 ∘C. What is the value of k at 60 ∘C if Ea = 135 kJ/mol ?

Express your answer using two significant figures.
Answer #1

Az = A e talRtz w [ Temperature dependence of rate u) - Resistent I ki = A Ealati - lii) Diriling eqn ci) by equ lil we get Awerke 174% 10% meetrit 333K 2.303 x 8.31435 mort (298 2.30348. 314 XA x 298% *333K/ login or ) = 74x10° x 35% lego (por ) - 2B) logo los hotel ) liven: kg = 2140x 1025 T = 25°c = (25+273) K 2 2988 K2=? Izabolca (60+273) K = 333K Ea = 135kg/mol = 135x